OpenStudy (anonymous):

9x^2-9(a+b)x+(2a^2+5ab+2b^2)=0 Find x

OpenStudy (ash2326):

We have \[9x^2-9(a+b)x+(2a^2+5ab+2b^2)=0 \] We know standard quadratic equation \[ax^2+bx+c=0\] we know the roots are given by quadratic formula \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] here \[a=9\] \[b=-9(a+b)\] and \[c=2a^2+5ab+2b^2\] let's find the roots using quadratic formula \[x=\frac{-9(a+b)\pm \sqrt{(9(a+b))^2-4(9)(2a^2+5ab+2b^2)}}{2\times 9}\] now let's expand the terms \[x=\frac{-9(a+b)\pm \sqrt{(81a^2+81b^2+162ab-72a^2-180ab-72b^2)}}{18}\] we get now \[x=\frac{-9(a+b)\pm \sqrt{(9a^2+9b^2-18ab)}}{18}\] let's take 9 out of the square root we get \[x=\frac{-9(a+b)\pm 3\sqrt{(a^2+b^2-2ab)}}{18}\] we know that \[a^2-2ab+b^2=(a-b)^2\] we get now \[x=\frac{-9(a+b)\pm 3(a-b)}{18}\] so we'll have \[x=\frac{-9(a+b)+ 3(a-b)}{18} , \frac{-9(a+b)- 3(a-b)}{18}\] Let's simplify \[x=\frac{-6a-12b}{18} , \frac{-12a-6b}{18}\] let's divide numerator and denominator by 6 we get \[x=\frac{-a-2b}{3} , \frac{-2a-b}{3}\]

OpenStudy (anonymous):

you explain step by step.

OpenStudy (anonymous):

x=−a−2b | −2a−b ------- | ------- 3 | 3